OpenStudy (anonymous):

given sin 4/5 in quadrant 1 and cos -12/13 in quadrant 2 find cos (a+b)

OpenStudy (anonymous):

someone please respond im going to get a major whipping if i dont

OpenStudy (anonymous):

given cos -24/25 in quadrant 2 and sin -5/13 in quadrant 3 find sin (a+b)

OpenStudy (anonymous):

a. 204/325 b. -36/325 c. -204/325 d. 36/325 these are the answers to the second question

OpenStudy (anonymous):

u mean (cos a =-24/25) and (sin b=-5/13)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sin(a+b)=sina cosb + cosa sinb

OpenStudy (anonymous):

we have cosa and sinb so we need to find sina and cosb

OpenStudy (anonymous):

we know that \[\cos^2x+\sin^2x=1\]

OpenStudy (anonymous):

yes please continue

OpenStudy (anonymous):

\[\sin^2a=1-\cos^2a=1-(-24/25)^2=(7/25)^2\]

OpenStudy (anonymous):

\[\sin a=\pm 7/25\] we choose sin a =7/25 because a is in quadrant 2

OpenStudy (anonymous):

so is the answer a? 204/325?

OpenStudy (anonymous):

just like what we did for sina \[\cos b=\pm12/13\] we choose cosb =-12/13 because b is in quadrant 3

OpenStudy (anonymous):

wait a sec

OpenStudy (anonymous):

?

OpenStudy (anonymous):

answer is d

OpenStudy (anonymous):

tan5π/6+tanπ/61−tan5π/6tanπ/6

OpenStudy (anonymous):

a. √3 b. 0 c. undefined d. 1

OpenStudy (callisto):

@ap331 Please close this question and post the question as a new one when you ask another one :)

OpenStudy (ganpat):

@ap331: its little confusing. can u write back again ??